# On balanced Hermitian threefolds with parallel Bismut torsion

### v1

@@ 5.1. The local structure
To see that $L$ is actually a foliation, it suffices to show that $[e_{3},\overline{e}_{3}]=0$. Since $T(e_{i},\overline{e}_{j})=0$, by the expression for $\theta$ we have
$[e_{3},\overline{e}_{3}]=\nabla^{c}_{e_{3}}\overline{e}_{3}-\nabla^{c}_{\overline{e}_{3}}e_{3}=\sum_{k}\big{(}\overline{\theta_{3k}(\overline{e}_{3})}\,\overline{e}_{k}-\theta_{3k}(\overline{e}_{3})\,e_{3}\big{)}=0.$
@@
Let $(M^{3},g)$ be a compact balanced BTP threefold of middle type. A local unitary frame $e$ on $M^{3}$ is called an admissble frame, if under $e$ the non-zero Chern torsion components are $T^{1}_{13}=-T^{2}_{23}=a>0$.
@@
Also, under an admissble frame $e$, the Bismut connection matrix is diagonal and
@@
(5.3) $R^{b}_{i\bar{j}k\bar{\ell}}=R^{b}_{k\bar{\ell}i\bar{j}},\ \ \ \ \ R^{b}_{i\bar{j}k\bar{\ell}}-R^{b}_{k\bar{j}i\bar{\ell}}=\sum_{s}\big{(}T^{j}_{ks}\overline{T^{i}_{\ell s}}+T^{\ell}_{is}\overline{T^{k}_{js}}-T^{j}_{is}\overline{T^{k}_{\ell s}}-T^{\ell}_{ks}\overline{T^{i}_{js}}-T^{s}_{ik}\overline{T^{s}_{j\ell}}\big{)}$
for any indices $i$, $j$, $k$, $\ell$. The first equality is part (2) of Proposition 2.6 in [35], and the second one is by Definition 2.3 and formula (2.4) in Proposition 2.5 of [35]. In particular, for our $(M^{3},g)$ under an admissble frame $e$, by the first equation of (5.3) and the fact that $\Theta^{b}_{3\ast}=\Theta^{b}_{12}=0$, we know that the only possibly non-zero Bismut curvature components are $R^{b}_{1\bar{1}1\bar{1}}$, $R^{b}_{2\bar{2}2\bar{2}}$, and $R^{b}_{1\bar{1}2\bar{2}}=R^{b}_{2\bar{2}1\bar{1}}$. By the second equality of (5.3), for $1\leq i\neq k\leq 3$, we get
@@
[…]
@@
Let $(M^{3},g)$ be a compact balanced BTP threefold of middle type. Then the Bismut curvature is determined by two real-valued smooth functions $s$ and $\sigma_{2}$ on $M^{3}$. If it is locally homogeneous, then both $s$ and $\sigma_{2}$ are constants. Conversely, if both $s$ and $\sigma_{2}$ are constants, then as observed in [24, Proposition 1.7], $(M^{3},g)$ is actually Bismut Ambrose-Singer (abbreviated as BAS, which means that the Bismut connection has parallel torsion and curvature), in particular it is locally homogeneous. For the convenience of readers we include the sketch of proof below.
Lemma 5.6 ([24]).
@@
[…]
Given any Hermitian metric $g_{0}$ on $M^{3}$ with Kähler form $\omega_{0}$, locally we have $\omega_{0}=\sqrt{-1}\sum_{i,j}h_{i\bar{j}}\varphi_{i\bar{j}}$, where $(h_{i\bar{j}})$ is positive definite. Let $f$ be the function on $M^{3}$ defined by ¯$\partial\overline{\partial}\Phi\wedge\omega_{0}=f\omega^{3}$, then by
¯$\partial\overline{\partial}\Phi\wedge\omega_{0}=2a^{2}\varphi_{1\bar{1}}\varphi_{2\bar{2}}\wedge\omega_{0}=-\frac{1}{3}a^{2}h_{3\bar{3}}\,\omega^{3},$
@@
Recall that the Fino-Vezzoni Conjecture ([15, 16]) states that any compact complex manifold admitting both a balanced and a pluriclosed metric must be Kählerian, while Streets-Tian Conjecture [29] states that any Hermitian-symplectic manifold must be Kählerian. Supplementing Corollary 1.19 of [35], we now have the following:
@@
For balanced (but non-Kähler) BTP threefolds, the Fano case is already Kählerian, while the middle type ones do not admit any pluriclosed metric by the above proposition, hence Fino-Vezzoni Conjecture holds, and the Streets-Tian Conjecture also holds as any Hermitian-symplectic metric is necessarily pluriclosed. We are only left with the Chern flat case. Streets-Tian Conjecture is known to be true for compact Chern flat manifolds in all dimensions by the work of Di Scala-Lauret-Vezzoni [14, Proposition 3.3], while in dimension $3$, both conjectures hold since any compact, non-Kähler Chern flat threefold $(M^{3},g)$ cannot admit any pluriclosed metric, because $\sqrt{-1}\partial\overline{\partial}\omega=\,^{t}\!\tau\wedge\overline{\tau}$, where $\tau$ is column vector of Chern torison $(2,0)$-forms. If $g_{0}$ is a pluriclosed metric on $M^{3}$, then one would have
@@
which is a contradiction. Note that for the non-balanced case, both conjectures are valid in complex dimension $3$ by [35, Corollary 1.19]. ∎
@@ 5.2. A Hermitian Lie algebra example
In [35, Proposition 1.10], we see characterization of BTP Hermitian structures on nilpotent Lie algebras. In particular, in complex dimension $3$ there is only one (up to scaling the metric by a constant multiple) balanced BTP complex nilmanifold $N^{3}$ with nilpotent $J$ in the sense of [13], whose structure equation is given by
@@
where $a>0$ is a constant, $\varphi$ is a unitary coframe, and as before we have abbreviated $\varphi_{i}\wedge\overline{\varphi}_{j}$ as $\varphi_{i\bar{j}}$ for convenience. In this subsection, we will drop the nilpotency requirement and figure out all Hermitian Lie algebras in complex dimension $3$ that are balanced BTP of middle type. It turns that there are exactly two families of such Lie algebras, $A_{s,t}$ parameterized by $(s,t)\in{\mathbb{R}}^{2}$, and $B_{z,t}$ parameterized by $(z,t)\in{\mathbb{C}}\times{\mathbb{R}}$, such that at the origin they are just $N^{3}$, namely, $A_{0,0}=B_{0,0}=N^{3}$. Note that the two families have an overlap: $A_{0,t}=B_{0,t}$ for any $t$, and here we have assumed that the Lie algebra is unimodular, which is a necessary condition for the corresponding Lie group to admit a compact quotient. Other than $N^{3}$, these $A_{s,t}$ or $B_{z,t}$ are not nilpotent, instead they all $3$-step solvable. With the exception of $A_{-t,t}$, all others are not of Calabi-Yau type, namely, there is no non-trivial invariant holomorphic $3$-form.
@@
$[x,y]-[Jx,Jy]+J[Jx,y]+J[x,y]=0,\ \ \ \ \ \forall\ x,y\in{\mathfrak{g}}.$
Let $g=\langle\cdot,\cdot\rangle$ be a metric (inner product) on ${\mathfrak{g}}$ compatible with $J$, that is, $\langle Jx,Jy\rangle=\langle x,y\rangle$, $\forall$ $x,y\in{\mathfrak{g}}$. We will call $({\mathfrak{g}},J,g)$ a Hermitian Lie algebra (or Lie algebra equipped with a Hermitian structure). It corresponds to Lie groups equipped with a left-invariant complex structure and a compatible left-invariant metric. Write ${\mathfrak{g}}^{1,0}=\{x-\sqrt{-1}Jx\mid x\in{\mathfrak{g}}\}$. A unitary frame $e$ of ${\mathfrak{g}}$ is a unitary basis of ${\mathfrak{g}}^{1,0}$. Its dual coframe is a basis $\varphi$ of the dual vector space $({\mathfrak{g}}^{1,0})^{\ast}$ satisfying $\varphi_{i}(e_{j})=\delta_{ij}$ for all $1\leq i,j\leq n$. Following [32] and [12], let us denote the structure constants by
[…]
@@
[…]
@@
¯$\overline{\partial}\alpha=x_{2}\overline{\partial}\varphi_{2}+x_{3}d\varphi_{3}=-x_{2}\varphi_{2}\overline{\beta}+x_{3}d\varphi_{3}=ax_{3}(\varphi_{2\bar{2}}-\varphi_{1\bar{1}}).$
By ¯$\overline{\partial}\alpha-\partial\overline{\alpha}=\lambda_{1}\varphi_{1\bar{1}}$, we get $\lambda_{1}=x_{3}+\overline{x}_{3}=0$. By the same way, we conclude that $y_{1}x_{2}=y_{1}x_{3}=0$ and $\lambda_{2}=y_{3}+\overline{y}_{3}=0$. In summary, we have proved that

### v2

@@ 5.1. The local structure
To see that $L$ is actually a foliation, it suffices to show that $e_{3},\overline{e}_{3}]=0$. Since $T(e_{i},\overline{e}_{j})=0$, by the expression for $\theta$ we have
$e_{3},\overline{e}_{3}]=\nabla^{c}_{e_{3}}\overline{e}_{3}-\nabla^{c}_{\overline{e}_{3}}e_{3}=\sum_{k}\big(\overline{\theta_{3k}(\overline{e}_{3})}\,\overline{e}_{k}-\theta_{3k}(\overline{e}_{3})\,e_{3}\big)=0.$
@@
Let $(M^{3},g)$ be a compact balanced BTP threefold of middle type. A local unitary frame $e$ on $M^{3}$ is called an admissible frame, if under $e$ the non-zero Chern torsion components are $T^{1}_{13}=-T^{2}_{23}=a>0$.
@@
Also, under an admissible frame $e$, the Bismut connection matrix is diagonal and
@@
(5.3) $R^{b}_{i\bar{j}k\bar{\ell}}=R^{b}_{k\bar{\ell}i\bar{j}},\ \ \ \ \ R^{b}_{i\bar{j}k\bar{\ell}}-R^{b}_{k\bar{j}i\bar{\ell}}=\sum_{s}\big(T^{j}_{ks}\overline{T^{i}_{\ell s}}+T^{\ell}_{is}\overline{T^{k}_{js}}-T^{j}_{is}\overline{T^{k}_{\ell s}}-T^{\ell}_{ks}\overline{T^{i}_{js}}-T^{s}_{ik}\overline{T^{s}_{j\ell}}\big)\ \ \$
for any indices $i$, $j$, $k$, $\ell$. The first equality is part (2) of Proposition 2.6 in [ZZ24+], and the second one is by Definition 2.3 and formula (2.4) in Proposition 2.5 of [ZZ24+]. In particular, for our $(M^{3},g)$ under an admissible frame $e$, by the first equation of (5.3) and the fact that $\Theta^{b}_{3\ast}=\Theta^{b}_{12}=0$, we know that the only possibly non-zero Bismut curvature components are $R^{b}_{1\bar{1}1\bar{1}}$, $R^{b}_{2\bar{2}2\bar{2}}$, and $R^{b}_{1\bar{1}2\bar{2}}=R^{b}_{2\bar{2}1\bar{1}}$. By the second equality of (5.3), for $1\leq i\neq k\leq 3$, we get
@@
$3Ric(\omega)\wedge\omega^{2}=s\,\omega^{3},\ \ \ \ \ \ 3(Ric(\omega))^{2}\wedge\omega=\sigma_{2}\,\omega^{3},$
@@
Let $(M^{3},g)$ be a compact balanced BTP threefold of middle type. Then the Bismut curvature is determined by two real-valued smooth functions $s$ and $\sigma_{2}$ on $M^{3}$. If it is locally homogeneous, then both $s$ and $\sigma_{2}$ are constants. Conversely, if both $s$ and $\sigma_{2}$ are constants, then as observed in [PZ25, Proposition 1.7], $(M^{3},g)$ is actually Bismut Ambrose-Singer (abbreviated as BAS, which means that the Bismut connection has parallel torsion and curvature), in particular it is locally homogeneous. For the convenience of readers we include the sketch of proof below.
Lemma 5.6 ([PZ25]).
@@
$\partial\overline{\partial}\Phi\ =\ -d\partial\Phi\ =\ -d\big(a(\varphi_{2\bar{2}}-\varphi_{1\bar{1}})\wedge\varphi_{3}\big)\ =\ -a(\varphi_{2\bar{2}}-\varphi_{1\bar{1}})\wedge d\varphi_{3}\ =\ 2a^{2}\varphi_{1\bar{1}}\varphi_{2\bar{2}}.$
Given any Hermitian metric $g_{0}$ on $M^{3}$ with Kähler form $\omega_{0}$, locally we have $\omega_{0}=\sqrt{-1}\sum_{i,j}h_{i\bar{j}}\varphi_{i\bar{j}}$, where $(h_{i\bar{j}})$ is positive definite. Let $f$ be the function on $M^{3}$ defined by $\partial\overline{\partial}\Phi\wedge\omega_{0}=f\omega^{3}$, then by
$\partial\overline{\partial}\Phi\wedge\omega_{0}=2a^{2}\varphi_{1\bar{1}}\varphi_{2\bar{2}}\wedge\omega_{0}=-\frac{1}{3}a^{2}h_{3\bar{3}}\,\omega^{3},$
@@
Recall that the Fino-Vezzoni Conjecture ([FV15, FV16]) states that any compact complex manifold admitting both a balanced and a pluriclosed metric must be Kählerian, while Streets-Tian Conjecture [ST10] states that any Hermitian-symplectic manifold must be Kählerian. Supplementing Corollary 1.19 of [ZZ24+], we now have the following:
@@
For balanced (but non-Kähler) BTP threefolds, the Fano case is already Kählerian, while the middle type ones do not admit any pluriclosed metric by the above proposition, hence Fino-Vezzoni Conjecture holds, and the Streets-Tian Conjecture also holds as any Hermitian-symplectic metric is necessarily pluriclosed. We are only left with the Chern flat case. Streets-Tian Conjecture is known to be true for compact Chern flat manifolds in all dimensions by the work of Di Scala-Lauret-Vezzoni [1, Proposition 3.3], while in dimension $3$, both conjectures hold since any compact, non-Kähler Chern flat threefold $(M^{3},g)$ cannot admit any pluriclosed metric, because $\sqrt{-1}\partial\overline{\partial}\omega=\,^{t}\!\tau\wedge\overline{\tau}$, where $\tau$ is column vector of Chern torison $(2,0)$-forms. If $g_{0}$ is a pluriclosed metric on $M^{3}$, then one would have
@@
which is a contradiction. Note that for the non-balanced case, both conjectures are valid in complex dimension $3$ by [ZZ24+, Corollary 1.19]. ∎
@@ 5.2. A Hermitian Lie algebra example
In [ZZ24+, Proposition 1.10], we have seen the characterization of BTP Hermitian structures on nilpotent Lie algebras. In particular, in complex dimension $3$ there is only one (up to scaling the metric by a constant multiple) balanced BTP Hermitian nilpotent Lie algebra $N^{3}$ with nilpotent $J$ in the sense of [CFGU00], whose structure equation is given by
@@
where $a>0$ is a constant, $\varphi$ is a unitary coframe, and as before we have abbreviated $\varphi_{i}\wedge\overline{\varphi}_{j}$ as $\varphi_{i\bar{j}}$ for convenience. In this subsection, we will drop the nilpotency requirement and figure out all Hermitian Lie algebras in complex dimension $3$ that are balanced BTP of middle type. It turns that there are exactly two families of such Lie algebras, $A_{s,t}$ parameterized by $(s,t)\in{\mathbb{R}}^{2}$, and $B_{z,t}$ parameterized by $(z,t)\in{\mathbb{C}}\times{\mathbb{R}}$, such that at the origin they are just $N^{3}$, namely, $A_{0,0}=B_{0,0}=N^{3}$. Note that the two families have an overlap: $A_{0,t}=B_{0,t}$ for any $t$, and here we have assumed that the Lie algebra is unimodular, which is a necessary condition for the corresponding Lie group to admit a compact quotient. Other than $N^{3}$, these $A_{s,t}$ or $B_{z,t}$ are not nilpotent, instead they are all $3$-step solvable, which means that the underlying real Lie algebra $\mathfrak{g}$ satisfies
$\mathfrak{g}^{\prime}=[\mathfrak{g},\mathfrak{g}]\neq 0,\quad\mathfrak{g}^{\prime\prime}=[\mathfrak{g}^{\prime},\mathfrak{g}^{\prime}]\neq 0,\quad\mathfrak{g}^{\prime\prime\prime}=[\mathfrak{g}^{\prime\prime},\mathfrak{g}^{\prime\prime}]=0.$
With the exception of $A_{-t,t}$, all others are not of Calabi-Yau type, namely, there is no non-trivial invariant holomorphic $3$-form.
@@
$x,y]-[Jx,Jy]+J[Jx,y]+J[x,y]=0,\ \ \ \ \ \forall\ x,y\in{\mathfrak{g}}.\ \ \ \ \$
Let $g=\langle\cdot,\cdot\rangle$ be a metric (inner product) on $\mathfrak{g}$} compatible with $J$, that is, $\langle Jx,Jy\rangle=\langle x,y\rangle$, $\forall$ $x,y\in{\mathfrak{g}}$. We will call $({\mathfrak{g}},J,g)$ a Hermitian Lie algebra (or Lie algebra equipped with a Hermitian structure). It corresponds to Lie groups equipped with a left-invariant complex structure and a compatible left-invariant metric. Write ${\mathfrak{g}}^{1,0}=\{x-\sqrt{-1}Jx\mid x\in{\mathfrak{g}}\}$. A unitary frame $e$ of $\mathfrak{g}$} is a unitary basis of ${\mathfrak{g}}^{1,0}$. Its dual coframe is a basis $\varphi$ of the dual vector space $({\mathfrak{g}}^{1,0})^{\ast}$ satisfying $\varphi_{i}(e_{j})=\delta_{ij}$ for all $1\leq i,j\leq n$. Following [VYZ19] and [CZ25], let us denote the structure constants by
$C^{j}_{ik}=\varphi_{j}([e_{i},e_{k}]),\ \ \ \ \ D^{j}_{ik}=\overline{\varphi}_{i}([\overline{e}_{j},e_{k}]),\ \ \ \ \ 1\leq i,j,k\leq n.$
@@
(5.7) $d\varphi_{i}=-\frac{1}{2}\sum_{j,k}C^{i}_{jk}\,\varphi_{j}\wedge\varphi_{k}-\sum_{j,k}\overline{D^{j}_{ik}}\,\varphi_{j}\wedge\overline{\varphi}_{k},\ \ \ \ \ \ \ 1\leq i\leq n,$
@@
$\overline{\partial}\alpha=x_{2}\overline{\partial}\varphi_{2}+x_{3}d\varphi_{3}=-x_{2}\varphi_{2}\overline{\beta}+x_{3}d\varphi_{3}=ax_{3}(\varphi_{2\bar{2}}-\varphi_{1\bar{1}}).$
By $\overline{\partial}\alpha-\partial\overline{\alpha}=\lambda_{1}\varphi_{1\bar{1}}$, we get $\lambda_{1}=x_{3}+\overline{x}_{3}=0$. By the same way, we conclude that $y_{1}x_{2}=y_{1}x_{3}=0$ and $\lambda_{2}=y_{3}+\overline{y}_{3}=0$. In summary, we have proved that